Find the locus of the point z for given condition :avg z-1/z+1=π/6

Grade:12th pass

1 Answers

Vikas TU
7 years ago
put z= x+iy
we get,
(x + iy – 1)/(x + iy + 1) = 1/root(3)
on solving we get,
x = (root3 +1)/(root3 -1)
and y = -1/(1 + root(3))
multyiplying both we get,
xy = 1/(root(3) – 1)
Hyeperbola is the locus probably.